OpenStudy (anonymous):

What is the pH of a solution that has [HCN]=0.25M and [NaCN]=0.50M ? (ka of HCN=4.9 x 10^-10)

OpenStudy (aaronq):

use the Henderson-Hasslebalch equation: \(pH=pKa+log\dfrac{[CN^-]}{[HCN]}\)

OpenStudy (anonymous):

@aaronq thank you so much ! For this problem first I need to find pKa ? or can I just plug in my given values ?

OpenStudy (aaronq):

no problem! yeah find the the pKa is just the -log of the ka; pKa=-log(Ka) then plug your values in.

OpenStudy (anonymous):

Got it thank you so much !

OpenStudy (aaronq):

good stuff, no problem !